In(x-1)+In6= In(3x)
Answer by ankor@dixie-net.com(22740)

In(x-1)+ In(6) = In(3x)
adding logs mean multiply so we can write this as
ln[6(x-1)] = ln(3x)
therefore
6(x-1) = 3x
6x - 6 = 3x
6x - 3x = +6
3x = 6
x = 2
:
:
replace x with 2, in the original problem:
ln(2-1) + ln(6) = ln(2*3)
of course the ln of 1 is 0, so it checks out